OpenStudy (anonymous):

Evaluate the iterated integrals: 3 1 sqrt(1-z^2) ∫ ∫ ∫ ze^y dxdzdy 0 0 0 Please explain step by step

OpenStudy (anonymous):

okay so first step: isolate the innermost integral (\[\int\limits_{0}^{\sqrt{1-z^2}}ze^ydx\]

OpenStudy (anonymous):

since you have dx at the end, you'll be integrating with respect to x. So, you can treat all other variables as constants

OpenStudy (anonymous):

are you with me so far?

OpenStudy (anonymous):

yes but there is no x so would it be xze^y?

OpenStudy (anonymous):

yep! then plug in the limits for x

OpenStudy (anonymous):

so i would get (\[\sqrt{1-z ^{2}}ze ^{y}\]

OpenStudy (anonymous):

yes now isolate the second integral

OpenStudy (anonymous):

could i sub u = 1- z^2 and get du = -2zdz and dz= du/-2z

OpenStudy (anonymous):

and end up getting ∫ uze^y (du/-2z)

OpenStudy (anonymous):

which would elnimate z and ive me -1/2∫ u^3/3 (e^y) evaluated from 0 to 1?

OpenStudy (anonymous):

good except you'll integrate u^(1/2)*e^y

OpenStudy (anonymous):

and substituting the limits youll get from 1 to 0 so multiply the integral by -1 and integrate from 0 to 1

OpenStudy (anonymous):

Thanks so much i got it!

OpenStudy (anonymous):

you're welcome!